# On computability of equilibrium states

### v2

@@ 2.3. Ruelle–Perron–Frobenius operators and Gibbs states
for all $n\in\mathbb{N}$ and $x,y\in X$ with $\rho(x,y)\leqslant\xi$. For each $n\in\mathbb{N}$, set $u_{n}\coloneqq\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})$. By (2.11) and (2.20), $\{u_{n}\}_{n\in\mathbb{N}}$ is a uniformly bounded sequence of equicontinuous functions on $X$. By the Arzelà–Ascoli Theorem, there exists a continuous function $u_{\varphi}$ and an increasing sequence $\{n_{i}\}_{i\in\mathbb{N}}$ such that $u_{n_{i}}\rightarrow u_{\varphi}$ uniformly on $X$ as $i\rightarrow+\infty$. Thus it follows from the definition of $\{u_{n}\}_{n\in\mathbb{N}}$, (2.9), (2.11), and (2.14) that $u_{\varphi}$ satisfies (2.15), (2.16), (2.17), and (2.18). Next, we demonstrate that $u_{\varphi}$ is the unique subsequential limit of $\{u_{n}\}_{n\in\mathbb{N}}$. Suppose that $v_{\varphi}$ is an arbitrary subsequential limit of $\{u_{n}\}_{n\in\mathbb{N}}$. Then $v_{\varphi}$ is also a continuous function satisfying (2.15), (2.16), (2.17), and (2.18). Set
$t\coloneqq\sup\{s\in\mathbb{R}:u_{\varphi}(x)-sv_{\varphi}(x)>0\text{ for all }x\in X\}.$
Then there exists $y\in X$ such that $u_{\varphi}(y)=tv_{\varphi}(y)$. Since $\mathcal{L}^{n}_{\overline{\varphi}}(u_{\varphi}-tv_{\varphi})(y)=u_{\varphi}(y)-tv_{\varphi}(y)=0$ for each $n\in\mathbb{N}$, we obtain that $u_{\varphi}(x)-tv_{\varphi}(x)=0$ for each $x\in\bigcup_{i\in\mathbb{N}}T^{-i}(y)$. Since $T$ is topologically transitive, the set $\bigcup_{i\in\mathbb{N}}T^{-i}(y)$ is dense in $X$. Together with the continuity of $u_{\varphi}-tv_{\varphi}$, $u_{\varphi}(x)-tv_{\varphi}(x)=0$ for each $x\in X$. Because $\int\!u_{\varphi}\,\mathrm{d}m=\int\!v_{\varphi}\,\mathrm{d}m=1$, we obtain $t=1$. This implies that $u_{\varphi}=v_{\varphi}$. Therefore $u_{n}\rightarrow u_{\varphi}$ uniformly on $X$ as $n\rightarrow+\infty$. By [PU10, Proposition 5.2.11], $m$ is a Gibbs state for $T$ and $\varphi$. According to [PU10, Proposition 5.1.1] and (2.19), $\mu$ is also a Gibbs state for $T$ and $\varphi$. Recall that for each $m_{0}\in\mathcal{P}(X)$, a linear operator $\mathcal{L}_{m_{0}}\colon L^{1}(m_{0})\rightarrow L^{1}(m_{0})$ is defined in [PU10, Chapter 5.2]. Here, for the Gibbs state $m$, $\mathcal{L}_{m}$ equals exactly $\mathcal{L}_{\overline{\varphi}}$. Then by [PU10, Proposition 5.2.2], it follows from (2.15) that $\mu$ is $T$-invariant. By [PU10, Corollary 5.2.14], one can see that $\mu$ is the unique $T$-invariant Gibbs state for $T$ and $\varphi$. ∎
@@ 3. Preliminaries on Computable Analysis
In this section, we recall some notions and results in Computable Analysis. The definitions we adopted in this section are consistent with [Weih00]. Consequently, it is convenient to think of the algorithms or machines mentioned below as Type-2 machines defined in [Weih00, Definition 2.1.1]. For more details, we refer the reader to [BBRY11, Section 3], [GHR10, Section 2], and [Weih00].
@@ 3.1. Algorithms and computability over the reals
    computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that for each $n\in\mathbb{N}$, $\bigl{\lvert}\frac{f(n)}{g(n)}-x\bigr{\rvert}<2^{-n}.$
@@
    lower-computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that $\bigl{\{}\frac{f(n)}{g(n)}\bigr{\}}_{n\in\mathbb{N}}$ is an increasing sequence and converges to $x$ as $n\rightarrow+\infty$.
  3. (iii)
    upper-computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that $\bigl{\{}\frac{f(n)}{g(n)}\bigr{\}}_{n\in\mathbb{N}}$ is a decreasing sequence and converges to $x$ as $n\rightarrow+\infty$.
@@ 3.2. Computable metric spaces
    there exists an algorithm which, on input $i,\,j,\,m\in\mathbb{N}$, outputs $y_{i,j,m}\in\mathbb{Q}$ satisfying $\lvert y_{i,j,m}-\rho(s_{i},s_{j})\rvert<2^{-m}$.
The points in $\mathcal{S}$ are said to be ideal. Using the canonical bijection between $\mathbb{N}^{3}$ and $\mathbb{N}$ (see Subsection 2.1), we fix the effective enumeration of ideal balls $\{B_{i}\}_{i\in\mathbb{N}}$ for the set $\{B(s_{i},j/k):i,\,j,\,k\in\mathbb{N}\}$ of balls with rational radii centered at points in $\mathcal{S}$. These balls are called the ideal balls in $(X,\,\rho,\,\mathcal{S})$.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. We say that a point $x\in X$ is computable if there exists a computable function $f\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\rho(s_{f(n)},x)<2^{-n}$ for each $n\in\mathbb{N}.$ Moreover, a sequence of points $\{x_{i}\}_{i\in\mathbb{N}}$ is said to be a uniformly computable sequence of points if there exists a computable function $f\colon\mathbb{N}^{2}\rightarrow\mathbb{N}$ such that $\rho(s_{f(n,m)},x_{m})<2^{-n}$ for all $n,\,m\in\mathbb{N}$.
As a remark, a finite sequence of computable points is always a uniformly computable sequence of points. Similarly, for other definitions of computable objects detailed below, we always have that a finite sequence of computable objects is always a uniformly computable sequence.
@@
As a consequence of Definitions 3.5 and 3.6, the canonical bijection between $\mathbb{N}^{2}$ and $\mathbb{N}$ (see Subsection 2.1) gives the following proposition.
@@
Before the definition of computable functions between computable metric spaces, we recall the definition of an oracle of a point in computable metric space.
@@
Given a computable metric space $(X,\,\rho,\,\mathcal{S})$, and a point $x\in X$, we say that a function $\varphi\colon\mathbb{N}\rightarrow\mathbb{N}$ is an oracle for $x\in X$ if $\rho(s_{\varphi(m)},x)<2^{-m}$ for each $m\in\mathbb{N}.$
@@
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces. Then a function $f\colon X\rightarrow X^{\prime}$ is computable if there exists an algorithm which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $n\in\mathbb{N}$ and an oracle $\varphi$ for $x$, outputs $m\in\mathbb{N}$ satisfying that $\rho^{\prime}(s^{\prime}_{m},f(x))<2^{-n}$. Moreover, a sequence $\{f_{i}\}_{i\in\mathbb{N}}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is called a uniformly computable sequence of functions if there exists an algorithm which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $i,\,n\in\mathbb{N}$, and an oracle $\varphi$ for $x$, outputs $m\in\mathbb{N}$ satisfying that $\rho^{\prime}(s^{\prime}_{m},f_{i}(x))<2^{-n}$.
For example, in [Weih00], Examples 4.3.3 and 4.3.13.5 give the computability of the exponential function $\exp\colon\mathbb{R}\rightarrow\mathbb{R}$ and the logarithmic function $\log\colon\mathbb{R}^{+}\rightarrow\mathbb{R}$, respectively. The following proposition is a classical result that describes a topological property of a computable function.
@@
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\{s_{i}^{\prime}\}_{i\in\mathbb{N}}$, and $\{B_{i}^{\prime}\}_{i\in\mathbb{N}}$ is the effective enumeration of ideal balls of $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$. If $f\colon X\rightarrow X^{\prime}$ is computable, then $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$, and preimages of lower-computable open sets are still lower-computable open sets.
Proof.
Assume that $f\colon X\rightarrow X^{\prime}$ is a computable function. First, we demonstrate that $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Fix an integer $i\in\mathbb{N}$. Now we compute $\{c_{j}\}_{j\in\mathbb{N}}$ and $\{r_{j}\}_{j\in\mathbb{N}}$ satisfying that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Since $\{B_{i}^{\prime}\}_{i\in\mathbb{N}}$ is an effective enumeration, we can compute $m\in\mathbb{N}$ and $r\in\mathbb{Q}$ satisfying that $B_{\rho^{\prime}}(s_{m}^{\prime},r)=B_{i}^{\prime}$. Then we say that a sequence $\{p_{i}\}_{i=1}^{q}$ of integers is admissible if $B_{\rho}\bigl{(}s_{p_{i+1}},2^{-i-1}\bigr{)}\subseteq B_{\rho}\bigl{(}s_{p_{i}},2^{-i}\bigr{)}$ for each $i\in\mathbb{N}$. Note that $(X,\,\rho,\,\mathcal{S})$ is a computable metric space. It follows from statement (iii) of Definition 3.3 that we can compute all the possible admissible sequences. We enumerate them by $\{P_{i}\}_{i\in\mathbb{N}}$ effectively. Moreover, for each admissible sequence $P=\{p_{i}\}_{i=1}^{q}$, the function $\varphi_{P}\colon\mathbb{N}\rightarrow\mathbb{N}$ given by
$\varphi_{P}(i)\coloneqq\begin{cases}p_{i}&\text{if }1\leqslant i\leqslant q;\\ p_{q}&\text{if }i\geqslant q+1\end{cases}$
is an oracle for $s_{p_{q}}$. Since $f$ is computable, there exists an algorithm $M_{f}$ which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $n\in\mathbb{N}$ and an oracle $\varphi$ for $x$, outputs $l\in\mathbb{N}$ satisfying that $\rho^{\prime}(s^{\prime}_{l},f(x))<2^{-n}$. Now we design the algorithm $M_{f}^{\prime}$ to compute $\{c_{j}\}_{j\in\mathbb{N}}$ and $\{r_{j}\}_{j\in\mathbb{N}}$ satisfying that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$ as follows. We enumerate $\mathbb{N}\times\mathbb{N}$ as $\{(i_{u},n_{u})\}_{u\in\mathbb{N}}$ effectively.
  1. (i)
    Set $u$ to be $1$ and $j$ to be $1$.
  2. (ii)
    We input the oracle $\varphi_{P_{i_{u}}}$ and $n_{u}$ in $M_{f}$ to obtain the output $m_{i_{u},n_{u}}$. Assume that $M_{f}$ only reads $\{\varphi_{P_{i_{u}}}(v):v\in[1,k_{i_{u},n_{u}}]\cap\mathbb{N}\}$ to obtain this output.
  3. (iii)
    We check whether $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}\subseteq B_{\rho^{\prime}}(s^{\prime}_{m},r)$. If so, then define $c_{j}\coloneqq\varphi_{P_{i_{u}}}(k_{i_{u},n_{u}})$ and $r_{j}\coloneqq 2^{-k_{i_{u},n_{u}}}$, and set $j$ to be $j+1$.
  4. (iv)
    Set $u$ to be $u+1$ and come back to Step (ii).
Now we verify that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Fix an integer $j\in\mathbb{N}$. Now we demonstrate that $f\bigl{(}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}\bigr{)}\subseteq B_{i}^{\prime}$. By Step (iii) of $M_{f}^{\prime}$, we have $c_{j}=\varphi_{P_{i_{u}}}(k_{i_{u},n_{u}})$ and $r_{j}=2^{-k_{i_{u},n_{u}}}$ for some $u\in\mathbb{N}$ with $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}\subseteq B_{\rho^{\prime}}(s^{\prime}_{m},r)$. For each $x\in B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$, there exists an oracle $\widetilde{\varphi}_{P_{i_{u}},x}$ for $x$ satisfying that $\widetilde{\varphi}_{P_{i_{u}},x}(v)=\varphi_{P_{i_{u}}}(v)$ for each integer $v\in[1,k_{i_{u},n_{u}}]$. If we input $\widetilde{\varphi}_{P_{i_{u}},x}$ and $n_{u}$ in $M_{f}$, then the output is still $m_{i_{u},n_{u}}$. Then it follows from the definition of $M_{f}$ and Step (ii) of $M_{f}^{\prime}$ that $f(x)\in B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}$. To sum up, we have $f\bigl{(}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}\bigr{)}\subseteq B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}\subseteq B_{\rho^{\prime}}(s^{\prime}_{m},r)=B_{i}^{\prime}$. Thus, $f^{-1}(B_{i}^{\prime})\supseteq\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Next, we establish that $f^{-1}(B_{i}^{\prime})\subseteq\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Fix an arbitrary point $x\in f^{-1}(B_{i}^{\prime})$, it suffices to check $x\in B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$ for some $j\in\mathbb{N}$. Since $\mathcal{S}$ is dense in $X$, there exists an oracle $\overline{\varphi}_{x}$ for $x$ satisfying that $\{\overline{\varphi}_{x}(v)\}_{v=1}^{q}$ is an admissible sequence for each $q\in\mathbb{N}$. Since $f(x)\in B_{i}^{\prime}$, there exists $n(x)\in\mathbb{N}$ satisfying that $B_{\rho^{\prime}}\bigl{(}f(x),2^{-n(x)+1}\bigr{)}\subseteq B_{i}^{\prime}$. We input $\overline{\varphi}_{x}$ and $n(x)$ in $M_{f}$ to obtain the output $m(x)$, then $\rho^{\prime}\bigl{(}s^{\prime}_{m(x)},f(x)\bigr{)}<2^{-n(x)}$, hence, $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m(x)},2^{-n(x)}\bigr{)}\subseteq B_{\rho^{\prime}}\bigl{(}f(x),2^{-n(x)+1}\bigr{)}\subseteq B_{i}^{\prime}$. Assume that $M_{f}$ only reads $\{\overline{\varphi}_{x}(v):v\in[1,k(x)]\cap\mathbb{N}\}$ to obtain this output. Note that $Q=\{\overline{\varphi}_{x}(v)\}_{v=1}^{k(x)}$ is an admissible sequence. Then in Step (ii) of $M_{f}^{\prime}$, we will input $\varphi_{Q}$ and $n(x)$ in $M_{f}$. It follows from the definition of $k(x)$ and $Q$ that this output is still $m(x)$. Since $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m(x)},2^{-n(x)}\bigr{)}\subseteq B_{i}^{\prime}$, it follows from the definition of $\varphi_{Q}$ that $c_{j(x)}=\varphi_{Q}(k(x))=\overline{\varphi}_{x}(k(x))$ and $r_{j(x)}=2^{-k(x)}$ for some $j(x)\in\mathbb{N}$. By the definition of $\overline{\varphi}_{x}$, we have $x\in B_{\rho}\bigl{(}s_{\varphi_{x}(k(x))},2^{-k(x)}\bigr{)}=B_{\rho}\bigl{(}s_{c_{j(x)}},r_{j(x)}\bigr{)}$. To sum up, one can see that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B(s_{c_{j}},r_{j})$. Hence, by Definition 3.6, $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Therefore, it follows from Definition 3.5 and Proposition 3.7 that preimages of lower-computable open sets are still lower-computable open sets. ∎
@@
Fix an effective enumeration $\{q_{n}\}_{n\in\mathbb{N}}$ of $\mathbb{Q}$. Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $C$ be a subset of $X$. A function $T\colon X\rightarrow\mathbb{R}$ is said to be upper-computable on $C$ if there is a uniformly computable sequence $\{U_{i}\}_{i\in\mathbb{N}}$ of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$ satisfying that
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $T\colon X\rightarrow X$ a function with $p=p(x)\coloneqq\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}<+\infty$ for each $x\in X$. Then the inverse $T^{-1}$ of $T$ is said to be computable if there exists an algorithm $\mathcal{A}$ such that for each $x\in X$ and each $n\in\mathbb{N}$, we can input an oracle $\varphi$ for $x$ in $\mathcal{A}$ to produce $\{y_{i}:1\leqslant i\leqslant p(x)\}$ that satisfies the following property: If $T^{-1}(x)=\{x_{i}:1\leqslant i\leqslant p(x)\}$, then for each $1\leqslant i\leqslant p(x)$, there exists $1\leqslant j\leqslant p(x)$ such that $\rho(x_{i},y_{j})<2^{-n}.$
It follows immediately from Definition 3.9 that the computable real-valued functions are closed under finitely many operations from the following list: addition, multiplication, division, scalar multiplication, $\max,$ and $\min$ (see for example, [Weih00, Corollary 4.3.4]). At the end of this subsection, we recall the definitions of a recursively compact set and a recursively precompact metric space introduced in [GHR10, Section 2].
@@
In a computable metric space $(X,\,\rho,\,\mathcal{S})$, a set $K\subseteq X$ is said to be recursively compact if it is compact and there is an algorithm which, on input $\{i_{j}\}_{j=1}^{p}\subseteq\mathbb{N}$ and $\{q_{j}\}_{j=1}^{p}\subseteq\mathbb{Q}^{+}$, halts if and only if $K\subseteq\bigcup_{j=1}^{p}B(s_{i_{j}},q_{j})$.
[GHR10, Proposition 1] gives some basic properties of recursively compact sets.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Then the following statements hold:
@@
    Assume that $K\subseteq X$ is a recursively compact set, and $U\subseteq X$ is a lower-computable open set. Then $K\setminus U$ is recursively compact.

### v3

@@ 2.3. Ruelle–Perron–Frobenius operators and Gibbs states
for all $n\in\mathbb{N}$ and $x,y\in X$ with $\rho(x,y)<\xi$ and $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\geqslant\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)$. For each $n\in\mathbb{N}$, set $u_{n}\coloneqq\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})$. By (2.12) and (2.21), $\{u_{n}\}_{n\in\mathbb{N}}$ is a uniformly bounded sequence of equicontinuous functions on $X$. By the Arzelà–Ascoli theorem (see e.g. [Fo99, Theorem 4.44]), there exists a continuous function $u_{\varphi}$ and an increasing sequence $\{n_{i}\}_{i\in\mathbb{N}}$ such that $u_{n_{i}}\rightarrow u_{\varphi}$ uniformly on $X$ as $i\rightarrow+\infty$. Thus it follows from the definition of $\{u_{n}\}_{n\in\mathbb{N}}$, (2.10), (2.12), and (2.15) that $u_{\varphi}$ satisfies (2.16), (2.17), (2.18), and (2.19).
Next, it remains to prove that $u_{\varphi}$ is the unique subsequential limit of $\{u_{n}\}_{n\in\mathbb{N}}$. This part of the proof follows verbatim the same as the proof of [Li18, Theorem 5.16].
By [PU10, Proposition 5.2.11], $m$ is a Gibbs state for $T$ and $\varphi$. According to [PU10, Proposition 5.1.1] and (2.20), $\mu$ is also a Gibbs state for $T$ and $\varphi$. Finally, we shall prove that $\mu$ is $T$-invariant. It suffices to show that $\int\!(g\circ T)\,\mathrm{d}\mu=\int\!g\,\mathrm{d}\mu$ for each $g\in C(X)$. Indeed, by (2.20), (2.8), and (2.16), we obtain that
$\int\!(g\circ T)\,\mathrm{d}\mu$ $=\int\!u_{\varphi}\cdot(g\circ T)\,\mathrm{d}m=\int\!u_{\varphi}\cdot(g\circ T)\,\mathrm{d}\bigl{(}\mathcal{L}^{*}_{\overline{\varphi}}(m)\bigr{)}$
$=\int\!\mathcal{L}_{\overline{\varphi}}(u_{\varphi})\cdot g\,\mathrm{d}m=\int\!u_{\varphi}\cdot g\,\mathrm{d}m=\int\!g\,\mathrm{d}\mu.$
Therefore, by [PU10, Corollary 5.2.14], $\mu$ is the unique $T$-invariant Gibbs state for $T$ and $\varphi$. ∎
@@ 3. Preliminaries on Computable Analysis
In this section, we recall some notions and results from Computable Analysis. The definitions we adopt in this section are consistent with those in [Weih00]. Consequently, it is convenient to think of the algorithms or machines mentioned below as Type-2 machines defined in [Weih00, Definition 2.1.1]. For more details, we refer the reader to [BBRY11, Section 3], [GHR11, Section 2], and [Weih00].
@@ 3.1. Algorithms and computability over the reals
    computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that for each $n\in\mathbb{N}$, $\bigl{\lvert}\frac{f(n)}{g(n)}-x\bigr{\rvert}<2^{-n};$
@@
    lower-computable (resp. upper-computable) if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that $\bigl{\{}\frac{f(n)}{g(n)}\bigr{\}}_{n\in\mathbb{N}}$ is an increasing (resp. decreasing) sequence and converges to $x$ as $n\rightarrow+\infty$.
@@ 3.2. Computable metric spaces
    there exists an algorithm which, on input $i,\,j,\,m\in\mathbb{N}$, outputs $y_{i,j,m}\in\mathbb{Q}$ satisfying
    $\lvert y_{i,j,m}-\rho(s_{i},s_{j})\rvert<2^{-m}.$
The points in $\mathcal{S}$ are said to be ideal. Due to the existence of a computable bijection between $\mathbb{N}^{3}$ and $\mathbb{N}$, there exists an effective enumeration $\{B_{l}\}_{l\in\mathbb{N}}$ of the set $\{B(s_{i},j/k):i,\,j,\,k\in\mathbb{N}\}$ of balls with rational radii centered at points in $\mathcal{S}$. These balls are called the ideal balls in $(X,\,\rho,\,\mathcal{S})$.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. We say that a point $x\in X$ is computable if there exists a computable function $f\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\rho\bigl{(}s_{f(n)},x\bigr{)}<2^{-n}$ for each $n\in\mathbb{N}.$ Moreover, a sequence of points $\{x_{i}\}_{i\in\mathbb{N}}$ is said to be a uniformly computable sequence of points if there exists a computable function $f\colon\mathbb{N}^{2}\rightarrow\mathbb{N}$ such that $\rho\bigl{(}s_{f(n,m)},x_{m}\bigr{)}<2^{-n}$ for all $n,\,m\in\mathbb{N}$.
Remark.
For each finite sequence $\{x_{i}\}_{i=1}^{n}$ of computable points, we see it as $\{x_{i}\}_{i\in\mathbb{N}}$, where $x_{j}=x_{n}$ for each integer $j\geqslant n$. Then by Definition 3.4, it is not hard to see that $\{x_{i}\}_{i\in\mathbb{N}}$ is a uniformly computable sequence. As a convention, we will say that a finite sequence $\{x_{i}\}_{i=1}^{n}$ of computable points is a uniformly computable sequence directly. Similarly, for other definitions of computable objects detailed below, we always have that a finite sequence of computable objects is always a uniformly computable sequence.
@@
By Definitions 3.53.6, and the existence of computable bijections between $\mathbb{N}^{2}$ and $\mathbb{N}$, we obtain the following result.
@@
Before the definition of computable functions between computable metric spaces, we recall the definition of an oracle of a point in a computable metric space.
@@
Given a computable metric space $(X,\,\rho,\,\mathcal{S})$ and a point $x\in X$, we say that a function $\varphi\colon\mathbb{N}\rightarrow\mathbb{N}$ is an oracle for $x\in X$ if $\rho\bigl{(}s_{\varphi(m)},x\bigr{)}<2^{-m}$ for each $m\in\mathbb{N}.$
@@
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces. Then a function $f\colon X\rightarrow X^{\prime}$ is computable if there exists an algorithm which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $n\in\mathbb{N}$ and an oracle $\varphi$ for $x$, outputs $m\in\mathbb{N}$ satisfying $\rho^{\prime}(s^{\prime}_{m},f(x))<2^{-n}$. Moreover, a sequence $\{f_{i}\}_{i\in\mathbb{N}}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is called a uniformly computable sequence of functions if there exists an algorithm which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $i,\,n\in\mathbb{N}$, and an oracle $\varphi$ for $x$, outputs $m\in\mathbb{N}$ satisfying $\rho^{\prime}(s^{\prime}_{m},f_{i}(x))<2^{-n}$.
For example, in [Weih00], Examples 4.3.3 and 4.3.13.5 give the computability of the exponential function $\exp\colon\mathbb{R}\rightarrow\mathbb{R}$ and the logarithmic function $\log\colon\mathbb{R}^{+}\rightarrow\mathbb{R}$, respectively. The following proposition is a classical result that describes a topological property of a computable function (see e.g. [BH21, Proposition 5.2.14]).
@@
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\bigl{\{}s_{i}^{\prime}\bigr{\}}_{i\in\mathbb{N}}$, and $\bigl{\{}B_{i}^{\prime}\bigr{\}}_{i\in\mathbb{N}}$ is the effective enumeration of ideal balls of $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$. If $f\colon X\rightarrow X^{\prime}$ is computable, then $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$, and preimages of lower-computable open sets are still lower-computable open sets.
@@
Fix an effective enumeration $\{q_{n}\}_{n\in\mathbb{N}}$ of $\mathbb{Q}$. Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $C$ be a subset of $X$. A function $T\colon X\rightarrow\mathbb{R}$ is said to be upper-computable on $C$ if there is a uniformly computable sequence $\{U_{i}\}_{i\in\mathbb{N}}$ of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$ satisfying
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $T\colon X\rightarrow X$ a function with $p=p(x)\coloneqq\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}<+\infty$ for each $x\in X$. Then the inverse $T^{-1}$ of $T$ is said to be computable if there exists an algorithm $\mathcal{A}$ such that for each $x\in X$ and each $n\in\mathbb{N}$, we can input an oracle $\varphi$ for $x$ in $\mathcal{A}$ to produce $\{y_{i}:1\leqslant i\leqslant p(x)\}$ with the following property:
If $T^{-1}(x)=\{x_{i}:1\leqslant i\leqslant p(x)\}$, then for each $1\leqslant i\leqslant p(x)$, there exists $1\leqslant j\leqslant p(x)$ such that $\rho(x_{i},y_{j})<2^{-n}.$
It follows immediately from Definition 3.9 that the computable real-valued functions are closed under finitely many operations from the following list: addition, multiplication, division, scalar multiplication, $\max,$ and $\min$ (see e.g. [Weih00, Corollary 4.3.4]).
At the end of this subsection, we recall the definitions of a recursively compact set and a recursively precompact metric space introduced in [GHR11, Section 2].
@@
In a computable metric space $(X,\,\rho,\,\mathcal{S})$, a set $K\subseteq X$ is said to be recursively compact if it is compact and there is an algorithm which, on input a sequence $\{i_{j}\}_{j=1}^{p}$ in $\mathbb{N}$ and a sequence $\{q_{j}\}_{j=1}^{p}$ in $\mathbb{Q}^{+}$, halts if and only if $K\subseteq\bigcup_{j=1}^{p}B\bigl{(}s_{i_{j}},q_{j}\bigr{)}$.
The following are some fundamental properties of recursively compact sets as discussed in [GHR11, Proposition 1].
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $K\subseteq X$ is a recursively compact set and $f\colon X\rightarrow\mathbb{R}$ is a computable function. Then the following statements are true:
@@
    $X\smallsetminus K$ is a lower-computable open set.
  3. (iii)
    If $U\subseteq X$ is a lower-computable open set, then $K\smallsetminus U$ is recursively compact.
  4. (iv)
    $\inf_{x\in K}f(x)$ is lower-computable and $\sup_{x\in K}f(x)$ is upper-computable.
  5. (v)
    If $K^{\prime}\subseteq X$ is a recursively compact set, then so is $K^{\prime}\cap K$.